a pharmacist wishes to prepare 30ml 0f a 20% alcohol solution. She has 30% solution and 10% percent solution in her stock. How much of each should she use in forming the desired mixture?
Answer by robertb(5830)

0.30x + 0.10(30 - x) = 0.20*30,
0.30x + 3 - 0.10x = 6,
0.20x = 3,
x = 15 ml, the amount 30% alcohol solution, and
30- x = 15 ml, the amount of 10% alcohol solution.